OpenStudy (anonymous):

The length of a table is 5 cm more than its width. Its area is 126 cm2. Find its dimensions. A. 4 cm by 9 cm B. 9 cm by 14 cm C. 14 cm by 19 cm D. 19 cm by 24 cm

Parth (parthkohli):

Let the length be x + 5 centimetres, and the width be x centimetres. \( \color{Black}{\Rightarrow x( x + 5) = 126}\) \( \color{Black}{\Rightarrow x^2 + 5x = 126 }\) \( \color{Black}{\Rightarrow x^2 + 5x - 126 = 0}\)

Parth (parthkohli):

Solve this quadratic equation.

OpenStudy (anonymous):

And also I know you should show how you solved the problem, but if you multiply each of the possible answers, Answer C is the only answer that will result in an area of 126\[cm ^{2}\]